# Type Ia supernova growth-rate measurement with LSST simulations: intrinsic scatter systematics

### v1

@@ Abstract
Measurement of the growth rate of structures ($f\sigma_{8}$) with Type Ia supernovae (SNe Ia) will improve our understanding of the nature of dark energy and enable tests of general relativity. In this paper, we generate simulations of the 10 year SN Ia dataset of the Rubin-LSST survey, including a correlated velocity field from a N-body simulation and realistic models of SNe Ia properties and their correlations with host-galaxy properties. We find, similar to SN Ia analyses that constrain the dark energy equation-of-state parameters $w_{0}w_{a}$, that constraints on $f\sigma_{8}$ can be biased depending on the intrinsic scatter of SNe Ia. While for the majority of intrinsic scatter models we recover $f\sigma_{8}$ with a precision of $\sim 13-14\%$, for the most realistic dust-based model, we find that the presence of non-Gaussianities in Hubble diagram residuals leads to a bias on $f\sigma_{8}$ of about $\sim-20\%$. When trying to correct for the dust-based intrinsic scatter, we find that the propagation of the uncertainty on the model parameters does not significantly increase the error on $f\sigma_{8}$. We also find that while the main component of the error budget of $f\sigma_{8}$ is the statistical uncertainty ($>75\%$ of the total error budget), the systematic error budget is dominated by the uncertainty on the damping parameter, $\sigma_{u}$, that gives an empirical description of the effect of redshift space distortions on the velocity power spectrum. Our results motivate a search for new methods to correct for the non-Gaussian distribution of the Hubble diagram residuals, as well as an improved modeling of the damping parameter.
software: numpy (C. R. Harris et al., 2020), matplotlib (J. D. Hunter, 2007), astropy (T. P. Robitaille et al., 2013; The Astropy Collaboration et al., 2018, 2022), scipy (P. Virtanen et al., 2020), pandas (The pandas development team, 2024)
@@ 1 Introduction
In the past decades, Type Ia supernovae (SNe Ia) have been used as powerful tools to probe the content of our universe and test the standard model of cosmology. In particular, they have been widely used in measurements of the Hubble constant, $H_{0}$ (e.g. W. L. Freedman et al. 2019; A. G. Riess et al. 2022; D. Scolnic et al. 2023; L. Galbany et al. 2023), and of the dark energy equation-of-state parameter, $w$ (D. Brout et al., 2022; DES Collaboration et al., 2024). The number of available SNe Ia for cosmology have grown from a few hundred in the 2010’s to a few thousand today with datasets consisting of multiple surveys (D. Scolnic et al., 2022; B. O. Sánchez et al., 2024). This number is expected to increase by several orders of magnitude, reaching $\sim O(100~{}000)$ SNe Ia in the coming years with a new generation of full-sky surveys such as the Zwicky Transient Facility (ZTF, S. Dhawan et al. 2021; M. Rigault et al. 2025) in the Northern Hemisphere and the Vera C. Rubin Observatory Legacy Survey of Space and Time [^1] http://www.lsst.org. (LSST) in the Southern Hemisphere. These datasets will enable the use of SNe Ia in constraining new cosmological parameters such as the product of the growth rate of structure $f$ with the amplitude of matter density fluctuations $\sigma_{8}$, commonly noted $f\sigma_{8}$. The growth rate of structure, $f\sigma_{8}$, quantifies the rate of evolution of matter overdensities and velocity perturbations. Hence, this parameter is sensitive not only to dark energy, but also to the law of gravity and can be used to test the nature of dark energy and general relativity (D. Huterer, 2023; R. J. Turner, 2024). The measurement of the growth rate of structure is part of the scientific goals of the LSST Dark Energy Science Collaboration [^2] http://lsstdesc.org. (DESC, LSST Dark Energy Science Collaboration 2012). In this work, we prepare for the future measurement of $f\sigma_{8}$ using LSST SN Ia data by investigating some of the key systematics that could impact the analysis.
@@
Other methods beyond SNe Ia have also been used to constrain $f\sigma_{8}$ and can be divided into two main categories: indirect and direct measurements. Indirect measurements constrain $f\sigma_{8}$ through the redshift space distortion (RSD) caused by the Doppler effect from PVs on the correlation function of galaxy clustering (N. Kaiser, 1987). The most recent measurements using this technique are those from the Dark Energy Spectroscopic Instrument (DESI) first data release, (DESI Collaboration et al., 2024) which is a spectroscopic survey mapping large scale structures over a 14 000 deg2 area of the sky. These indirect measurements require a high number of galaxies and are more precise at high redshifts where there is more volume. Measurements of $f\sigma_{8}$ through RSD of galaxy clustering are expected to give a percent-level precision at the end of the DESI survey (DESI Collaboration et al., 2016).
@@
On the other hand, direct measurements of $f\sigma_{8}$ use estimation of the PVs and apply summary statistics to them in order to extract cosmological constraints. This type of measurement requires, in addition to spectroscopic redshifts, estimation of distances from a photometric survey in order to break the degeneracy between $z_{\mathrm{cos}}$ and $z_{p}$ in Eq. 1 and infer the galaxy PVs. This direct method is most efficient at low-$z$ where the amplitude of $z_{p}$ is non-negligible compared to the cosmological redshift $z_{\mathrm{cos}}$. Most of the existing direct constraints for $f\sigma_{8}$ use galaxies as distance indicators through the Tully-Fisher (TF, R. B. Tully & J. R. Fisher 1977) and Fundamental Plane (FP, S. Djorgovski & M. Davis 1987) relationships. After the estimation of PVs, several summary statistics can be applied to constrain $f\sigma_{8}$. These summary statistics include the compressed two-point correlation function (A. Nusser, 2017; F. Qin et al., 2019; R. J. Turner et al., 2022); the comparison between estimated velocities and velocities obtained from density field reconstruction (J. Carrick et al., 2015; S. S. Boruah et al., 2020; K. Said et al., 2020); the forward modelling reconstruction of the velocity field (S. S. Boruah et al., 2021; A. Valade et al., 2022; J. Prideaux-Ghee et al., 2022); and the maximum likelihood method (A. Abate & P. Erdogdu, 2009; A. Johnson et al., 2014; C. Howlett et al., 2017b). More recently, this has also been extended to cross-correlation analyses with PVs and density clustering data (C. Adams & C. Blake, 2020; Y. Lai et al., 2023).
@@
In this work, we focus on the direct constraint of $f\sigma_{8}$ with the maximum likelihood method using SNe Ia as distance indicators. Type Ia supernovae are better distance indicators than galaxies, with a relative precision of about $\sim$ 6% compared to $\sim 20\%$ for the distances inferred with TF and FP relationships. However, currently available low-$z$ SN Ia datasets are too small and not homogeneous enough to allow a SNe Ia only measurement of $f\sigma_{8}$, and most SN Ia cosmological analyses have instead treated PVs as a systematic (T. M. Davis et al., 2011; E. R. Peterson et al., 2022; A. Carr et al., 2022; B. Carreres et al., 2024; E. R. Peterson et al., 2025; A. M. Hollinger & M. J. Hudson, 2025; E. Tsaprazi & A. F. Heavens, 2025). A few analyses have made use of SN Ia data in combination with other probes to constrain $f\sigma_{8}$ (J. Carrick et al., 2015; D. Huterer et al., 2017; S. S. Boruah et al., 2021), but their methods were not extensively tested with simulations. In the context of the ZTF and LSST surveys, multiple Fisher forecast studies have been done and predict a precision for $f\sigma_{8}$ of a few percent with SNe Ia only and in combination with density obtained from spectroscopic surveys (J. Koda et al., 2014; C. Howlett et al., 2017a; R. Graziani et al., 2020; A. G. Kim & E. V. Linder, 2020). The first investigation of a SNe Ia only measurement of $f\sigma_{8}$ with simulations is presented in B. Carreres et al. (2023) (hereafter C23) within the context of the ZTF survey. This analysis showed that the 6 years of the ZTF survey can provide a constraint on $f\sigma_{8}$ with an error of 20% using a spectroscopically classified and complete ($z<0.06$) SNe Ia sample. However, C23 mainly focused on the selection function bias systematics, and additional systematics were neglected and left for future work. For instance, the simulations used in C23 are based on a dark matter halo $N$-body simulation and do not attempt to model correlations between SNe Ia and their host galaxies. Furthermore, the contamination of the SN Ia sample by core-collapse SNe (which will be addressed in Rosselli et al. in prep.) and the color dependence of the SN Ia intrinsic scatter were not taken into account. The latter, in particular, was recently found to be the most important systematic in the measurement of the dark energy equation of state parameter from the latest analysis of the Dark Energy Survey (DES) SN Ia data (M. Vincenzi et al., 2024, hereafter V24). The measurement of $f\sigma_{8}$ could be sensitive to correlations between color and velocity estimation or to possible non-Gaussianity introduced by the intrinsic scatter. To correct for potential biases that could arise from selection or color-dependent scattering while building the SN Ia hubble diagram, most recent SN Ia cosmological analyses (D. Brout et al., 2022; DES Collaboration et al., 2024) have relied on the BEAMS (M. Kunz et al., 2007) with Bias Corrections (BBC) framework (R. Kessler & D. Scolnic, 2017). In the BBC framework, large realistic simulations are used to compute a correction function of the biases of the Hubble diagram. This method has not yet been tested in the context of the $f\sigma_{8}$ measurement.
@@ 2 LSST Simulations
In the previous work of C23, simulations were run using a dark matter halo catalog as SN Ia hosts and no direct correlations were introduced between SNe Ia and their host galaxies. This could be problematic if such correlation introduces bias in the $f\sigma_{8}$ analysis, for instance, SN Ia host galaxies are known to be more massive and hence could have a biased velocity distribution. On a smaller scale, more massive galaxies are also brighter, increasing the noise on SN Ia flux measurement and thus on the estimated distances, which could also bias the velocity distribution. In this work we aim to simulate the peculiar velocity field at the galaxy level along with SNe Ia-host correlations.
@@ 2.1 Peculiar velocity field and host galaxy simulation
To produce the peculiar velocity field, the simulations of LSST SNe Ia are generated on top of the Uchuu UniverseMachine simulated galaxy catalog (T. Ishiyama et al., 2021; H. Aung et al., 2023). The Uchuu simulation consists of $12800^{3}$ dark matter particles of mass $m_{p}=3.27\times 10^{8}M_{\odot}$ in a $L_{\mathrm{Uchuu}}=2\ \mathrm{Gpc}~{}h^{-1}$ side length box. The simulation is initialized at a redshift of $z=127$ using the Planck15 results (Planck Collaboration et al., 2016) as the fiducial cosmology; the cosmological parameters are summarized in Table 1. As described in H. Aung et al. (2023), the Uchuu UniverseMachine galaxy catalog was constructed from the Uchuu $N$-body simulation with the UniverseMachine algorithm (P. Behroozi et al., 2019), which aims to forward model the relationship between halos and galaxy properties in order to match observed data. We make use of the various galaxy properties present in this catalog, such as stellar mass and star-formation rate (SFR), to correlate host and SN Ia properties. The Uchuu UniverseMachine data are available as 50 snapshots at different redshifts ranging from $z=14$ to $z=0$. In this work we use the lowest redshift snapshot at $z=0$.
We subdivide the $(2\ \mathrm{Gpc}~{}h^{-1})^{3}$ box into 8 non-overlapping sub-boxes of side length $L=1\ \mathrm{Gpc}~{}h^{-1}$ from which we produce 8 realizations of the 10 year LSST survey. We place the coordinate origin at the center of each box and convert distances to redshifts by numerically inverting the comoving distance-redshift relationship within the fiducial cosmology. The volume of each box corresponds to an effective redshift range of $z\in[0,0.175]$. This redshift range is sufficient to study the LSST SN Ia PV survey, as the noise of PV measurements increase with redshift.
In order to simulate the noise in SN Ia flux measurements due to the luminosity of the host galaxies, the magnitude and Sérsic profile parameters of the galaxies are interpolated in the $(\log M_{\mathrm{stellar}}/M_{\odot},\log\mathrm{SFR})$ parameter space from the galaxies produced for the OpenUniverse LSST-Roman simulations (OpenUniverse et al., 2025). The Sérsic parameters are also used to generate each SN Ia position relative to its host galaxy with a probability that depends on its profile.
As it was done in E. R. Peterson et al. (2025), the SN Ia hosts are randomly drawn according to the mass dependent SNe Ia rate described in Sect. 4 of P. Wiseman et al. (2021). In Fig. 1 we show the distribution of the masses of galaxies in the Uchuu UniverseMachine catalog compared to the distribution of host masses obtained in our simulation after applying the mass dependent random draw.
@@
Table 1: Fiducial cosmology parameters used in the simulation. These parameters are those from Planck Collaboration et al. (2016).
@@ 2.2 SN light-curve simulation and fitting
[…]
@@
To simulate a realistic number of SNe Ia  we use the SN Ia rate value from C. Frohmaier et al. (2019), which is $r_{v,\mathrm{F19}}=2.27\times(1+z)^{1.7}\times 10^{-5}$ SN Ia/Mpc/year and rescale it to our fiducial value of $h$ using
@@
The simulated photometry is obtained by the integration of a spectral template of each SN Ia. In this work, we use the spectral template from the SALT3 model (J. Guy et al., 2007; W. D. Kenworthy et al., 2021) as trained in V24. The SALT3 model gives a spectral template for SNe Ia that depends on three parameters: the magnitude of the SN Ia $m_{b}$, the stretch $x_{1}$, and the color $c$. The absolute magnitude of each SN Ia is defined by following the Tripp equation (R. Tripp, 1998) as
@@
where $\mu(\boldsymbol{z_{\mathrm{cos}}})$ is the distance moduli at the redshift of the SN Ia host and the last term is the relativistic beaming from peculiar velocity (T. M. Davis et al., 2011). In Eqs. 3 and 4 we put in bold the quantities that differ for each SN Ia.
We generate simulations using four different SN Ia intrinsic scatter models. The first is the least realistic and is used for comparison: coherent (achromatic) Gaussian scatter with a standard deviation of $\sigma_{M}=0.12$. We refer to this simulation as $S_{\mathrm{COH}}$. The next two are models that have been used historically in SN Ia analyses: the G10 model (J. Guy et al., 2010) where 70% of the luminosity scatter is due to achromatic variations and 30% to chromatic ones, and the C11 model (N. Chotard et al., 2011) where 25% of the scattering is achromatic and 70% is chromatic. These two models directly add a wavelength-dependent scatter to the spectral model provided by SALT3. We refer to these two simulations as $S_{\mathrm{G10}}$ and $S_{\mathrm{C11}}$, respectively. Recent studies, however, have disfavored these two models (V24). A more realistic model is presented in D. Brout & D. Scolnic (2021) (hereafter referred as BS21), which proposes that the scattering is caused by variation in dust extinction parameters of SN Ia host galaxies. We simulate the BS21 model using the dust parameters fitted in B. Popovic et al. (2023) (hereafter referred as P23). We refer to this simulation as $S_{\mathrm{P23}}$.
@@
with a value for $\gamma$ of 0.05 mag. The distribution of $x_{1}$ and $c$ are also chosen to be identical for these models. We use the parent populations $\mathcal{P}(x_{1}|M_{\mathrm{host}})$ and $\mathcal{P}(c|M_{\mathrm{host}})$ derived in Table 4 and 8 of B. Popovic et al. (2021) (hereafter referred as P21) which aim to model the correlations between SN Ia parameters and host masses.
In the BS21 model, the Tripp relation of Eq. 3 is slightly modified such as
@@
where $\alpha_{\mathrm{fid}}$ is again fixed at a value of 0.15 and $\mathcal{P}(x_{1}|M_{\mathrm{host}})$ is described by the same parent population as for the $S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$ and $S_{\mathrm{C11}}$ simulations. The BS21 model assumes that the SN color can be divided into two components. The first corresponds to an intrinsic color variation, $\beta c$, where $\beta$ is a random variable following a Gaussian law $\mathcal{P}(\beta)\sim\mathcal{N}(\mu_{\beta},\sigma_{\beta})$ of parameters
@@
$\tau_{E_{\mathrm{dust}}}=\left\{\begin{matrix}\left\{\begin{matrix}0.11&\text{if }\log M/M_{\odot}>10\\ 0.14&\text{ else}\end{matrix}\right.&\ \text{ if }z<0.1\\[8.61108pt] \left\{\begin{matrix}0.15&\text{ if }\log M/M_{\odot}>10\\ 0.12&\text{ else}\end{matrix}\right.&\ \text{ else }\end{matrix}\right..$ (10)
These values were derived using the method presented in P23 and are presented in Table 3 of V24. We summarize the parameters for our four intrinsic scatter models in Table 2.

### v2

@@ Abstract
Measurement of the growth rate of structures ($f\sigma_{8}$) with Type Ia supernovae (SNe Ia) will improve our understanding of the nature of dark energy and enable tests of general relativity. In this paper, we generate simulations of the 10 year SN Ia dataset of the Rubin-LSST survey, including a correlated velocity field from a N-body simulation and realistic models of SNe Ia properties and their correlations with host-galaxy properties. We find, similar to SN Ia analyses that constrain the dark energy equation-of-state parameters $w_{0}w_{a}$, that constraints on $f\sigma_{8}$ can be biased depending on the intrinsic scatter of SNe Ia. While for the majority of intrinsic scatter models we recover $f\sigma_{8}$ with a precision of $\sim$13–14%, for the most realistic dust-based model, we find that the presence of non-Gaussianities in Hubble diagram residuals leads to a bias on $f\sigma_{8}$ of about $\sim-20\%$. When trying to correct for the dust-based intrinsic scatter, we find that the propagation of the uncertainty on the model parameters does not significantly increase the error on $f\sigma_{8}$. We also find that while the main component of the error budget of $f\sigma_{8}$ is the statistical uncertainty ($75\%$ of the total error budget), the systematic error budget is dominated by the uncertainty on the damping parameter, $\sigma_{u}$, that gives an empirical description of the effect of redshift space distortions on the velocity power spectrum. Our results motivate a search for new methods to correct for the non-Gaussian distribution of the Hubble diagram residuals, as well as an improved modeling of the damping parameter.
software: numpy (harris_array_2020), matplotlib (Hunter:2007), astropy (robitaille_astropy_2013; the_astropy_collaboration_astropy_2018; the_astropy_collaboration_astropy_2022), scipy (virtanen_scipy_2020), pandas (the_pandas_development_team_pandas-devpandas_2024)
@@ 1 Introduction
In the past decades, Type Ia supernovae (SNe Ia) have been used as powerful tools to probe the content of our universe and test the standard model of cosmology. In particular, they have been widely used in measurements of the Hubble constant, $H_{0}$ (e.g. freedman_carnegie-chicago_2019; riess_comprehensive_2022; scolnic_cats_2023; galbany_updated_2023), and of the dark energy equation-of-state parameter, $w$ (brout_pantheon_2022; des_collaboration_dark_2024). The number of available SNe Ia for cosmology have grown from a few hundred in the 2010’s to a few thousand today with datasets consisting of multiple surveys (scolnic_pantheon_2022; sanchez_dark_2024). This number is expected to increase by several orders of magnitude, reaching $\sim O(100~000)$ SNe Ia in the coming years with a new generation of full-sky surveys such as the Zwicky Transient Facility (ZTF, dhawan_zwicky_2021; rigault_ztf_2025-1) in the Northern Hemisphere and the Vera C. Rubin Observatory Legacy Survey of Space and Time [^1] http://www.lsst.org. (LSST) in the Southern Hemisphere. These datasets will enable the use of SNe Ia in constraining new cosmological parameters such as the product of the growth rate of structure $f$ with the amplitude of matter density fluctuations $\sigma_{8}$, commonly noted $f\sigma_{8}$. The growth rate of structure, $f\sigma_{8}$, quantifies the rate of evolution of matter overdensities and velocity perturbations. Hence, this parameter is sensitive not only to dark energy, but also to the law of gravity and can be used to test the nature of dark energy and general relativity (huterer_growth_2023; turner_cosmology_2024). The measurement of the growth rate of structure is part of the scientific goals of the LSST Dark Energy Science Collaboration [^2] http://lsstdesc.org. (DESC, lsst_dark_energy_science_collaboration_large_2012). In this work, we prepare for the future measurement of $f\sigma_{8}$ using LSST SN Ia data by investigating some of the key systematics that could impact the analysis.
@@
Other methods beyond SNe Ia have also been used to constrain $f\sigma_{8}$ and can be divided into two main categories: indirect and direct measurements. Indirect measurements constrain $f\sigma_{8}$ through the redshift space distortion (RSD) caused by the Doppler effect from PVs on the correlation function of galaxy clustering (kaiser_clustering_1987). The most recent measurements using this technique are those from the Dark Energy Spectroscopic Instrument (DESI) second data release (desi_collaboration_desi_2025), which is a spectroscopic survey mapping large-scale structures over a 14 000 deg2 area of the sky. These indirect measurements require a high number of galaxies and are more precise at high redshifts where there is more volume. Measurements of $f\sigma_{8}$ through RSD of galaxy clustering are expected to give a percent-level precision at the end of the DESI survey (desi_collaboration_desi_2016).
@@
On the other hand, direct measurements of $f\sigma_{8}$ use PV estimations and apply summary statistics to them to extract cosmological constraints. This type of measurement requires, in addition to spectroscopic redshifts, estimation of distances from a photometric survey in order to break the degeneracy between $z_{\mathrm{cos}}$ and $z_{p}$ in Eq. 1 and infer the galaxy PVs. This direct method is most efficient at low-$z$ where the amplitude of $z_{p}$ is non-negligible compared to the cosmological redshift $z_{\mathrm{cos}}$. Most of the existing direct constraints for $f\sigma_{8}$ use galaxies as distance indicators through the Tully-Fisher (TF, tully_new_1977) and Fundamental Plane (FP, djorgovski_fundamental_1987) relationships. After the estimation of PVs, several summary statistics can be applied to constrain $f\sigma_{8}$. These summary statistics include the compressed two-point correlation function (nusser_velocitydensity_2017; qin_redshift_2019; turner_local_2022); the comparison between estimated velocities and velocities obtained from density field reconstruction (carrick_cosmological_2015; boruah_cosmic_2020; said_joint_2020); the forward modelling reconstruction of the velocity field (boruah_reconstructing_2021; valade_hamiltonian_2022; prideaux-ghee_field-based_2022); and the maximum likelihood method (abate_peculiar_2009; johnson_6df_2014; howlett_2mtf_2017). More recently, this has also been extended to cross-correlation analyses with PVs and density clustering data (adams_joint_2020; lai_using_2023).
@@
In this work, we focus on the direct constraint of $f\sigma_{8}$ with the maximum likelihood method using SNe Ia as distance indicators. Type Ia supernovae are better distance indicators than galaxies, with a relative precision of about $\sim$ 6% compared to $\sim 20\%$ for the distances inferred with TF and FP relationships. However, currently available low-$z$ SN Ia datasets are too small and not homogeneous enough to allow a SNe Ia only measurement of $f\sigma_{8}$, and most SN Ia cosmological analyses have instead treated PVs as a systematic (davis_effect_2011; peterson_pantheon_2022; carr_pantheon_2022; carreres_ztf_2025; peterson_improving_2025; hollinger_uncertainties_2025; tsaprazi_field-level_2025). A few analyses have made use of SN Ia data in combination with other probes to constrain $f\sigma_{8}$ (carrick_cosmological_2015; huterer_testing_2017; boruah_reconstructing_2021), but their methods were not extensively tested with simulations. In the context of the ZTF and LSST surveys, multiple Fisher forecast studies have been done and predict a precision for $f\sigma_{8}$ of a few percent with SNe Ia only and in combination with density obtained from spectroscopic surveys (koda_are_2014; howlett_cosmological_2017; graziani_peculiar_2020; kim_complementarity_2020). The first investigation of a SNe Ia only measurement of $f\sigma_{8}$ with simulations is presented in carreres_growth-rate_2023 (hereafter C23) within the context of the ZTF survey. This analysis showed that the 6 years of the ZTF survey can provide a constraint on $f\sigma_{8}$ with an error of 20% using a spectroscopically classified and complete ($z<0.06$) SNe Ia sample. However, carreres_growth-rate_2023 mainly focused on the selection function bias systematics, and additional systematics were neglected and left for future work. For instance, the simulations used in carreres_growth-rate_2023 are based on a dark matter halo $N$-body simulation and do not attempt to model correlations between SNe Ia and their host galaxies. Furthermore, the contamination of the SN Ia sample by core-collapse SNe (which will be addressed in rosselli_forecast_2025) and the color dependence of the SN Ia intrinsic scatter were not taken into account. The latter, in particular, was recently found to be the most important systematic in the measurement of the dark energy equation of state parameter from the latest analysis of the Dark Energy Survey (DES) SN Ia data (vincenzi_dark_2024, hereafter V24). The measurement of $f\sigma_{8}$ could be sensitive to correlations between color and velocity estimation or to possible non-Gaussianity introduced by the intrinsic scatter. To correct for potential biases that could arise from selection or color-dependent scattering while building the SN Ia Hubble diagram, most recent SN Ia cosmological analyses (brout_pantheon_2022; des_collaboration_dark_2024) have relied on the BEAMS (kunz_bayesian_2007) with Bias Corrections (BBC) framework (kessler_correcting_2017). In the BBC framework, large realistic simulations are used to compute a correction function of the biases of the Hubble diagram. This method has not yet been tested in the context of the $f\sigma_{8}$ measurement.
@@ 2 LSST Simulations
In the previous work of carreres_growth-rate_2023, simulations were run using a dark matter halo catalog as SN Ia hosts and no direct correlations were introduced between SNe Ia and their host galaxies. This could be problematic if such correlation introduces bias in the $f\sigma_{8}$ analysis, for instance, SN Ia host galaxies are known to be more massive and hence could have a biased velocity distribution. On a smaller scale, more massive galaxies are also brighter, increasing the noise on SN Ia flux measurement and thus on the estimated distances, which could also bias the velocity distribution. In this work we aim to simulate the peculiar velocity field at the galaxy level along with SNe Ia-host correlations.
@@ 2.1 Peculiar velocity field and host galaxy simulation
To produce the peculiar velocity field, the simulations of LSST SNe Ia are generated on top of the Uchuu UniverseMachine simulated galaxy catalog (ishiyama_uchuu_2021; aung_uchuu-universe_2023). The Uchuu simulation consists of $12800^{3}$ dark matter particles of mass $m_{p}=3.27\times 10^{8}M_{\odot}$ in a $L_{\mathrm{Uchuu}}=2\ \mathrm{Gpc}~h^{-1}$ side length box. The simulation is initialized at a redshift of $z=127$ using the Planck15 results (planck_collaboration_planck_2016) as the fiducial cosmology; the cosmological parameters are summarized in Table 1. As described in aung_uchuu-universe_2023, the Uchuu UniverseMachine galaxy catalog was constructed from the Uchuu $N$-body simulation with the UniverseMachine algorithm (behroozi_universemachine_2019), which aims to forward model the relationship between halos and galaxy properties in order to match observed data. We make use of the various galaxy properties present in this catalog, such as stellar mass and star-formation rate (SFR), to correlate host and SN Ia properties. The Uchuu UniverseMachine data are available as 50 snapshots at different redshifts ranging from $z=14$ to $z=0$. In this work we use the lowest redshift snapshot at $z=0$.
[…]
In order to simulate the noise in SN Ia flux measurements due to the luminosity of the host galaxies, the magnitude and Sérsic profile parameters of the galaxies are interpolated in the $(\log M_{\mathrm{stellar}}/M_{\odot},\log\mathrm{SFR})$ parameter space from the galaxies produced for the OpenUniverse LSST-Roman simulations (openuniverse_openuniverse2024_2025). The Sérsic parameters are also used to generate each SN Ia position relative to its host galaxy with a probability that depends on its profile.
As it was done in peterson_improving_2025, the SN Ia hosts are randomly drawn according to the mass-dependent SNe Ia rate described in Sect. 4 of wiseman_rates_2021. In Fig. 1 we show the distribution of the masses of galaxies in the Uchuu UniverseMachine catalog compared to the distribution of host masses obtained in our simulation after applying the mass dependent random draw.
@@
Table 1: Fiducial cosmology parameters used in the simulation. These parameters are those from planck_collaboration_planck_2016.
@@ 2.2 SN light-curve simulation and fitting
The LSST survey will use the Simonyi Survey Telescope at Rubin Observatory that collects light with a 8.4 meter primary mirror. The Rubin Observatory LSST Camera has a $\sim$ 9.4 deg2 field-of-view that will allow the LSST survey to observe the entire Southern Hemisphere sky multiple times across 10 years. We generate and analyze SN Ia data-like light-curves with the SNANA  [^3] https://github.com/RickKessler/SNANA. software library (kessler_snana_2009) run through the Pippin  [^4] https://github.com/dessn/Pippin. pipeline (hinton_pippin_2020). We simulate SN Ia data equivalent to 10 years of the LSST survey. To produce a realistic simulation of the LSST survey, SNANA requires two main inputs: a SIMLIB file that contains the observation characteristics and a HOSTLIB file that includes all the potential SN Ia host galaxies. These files are produced using the DESC software OpSimSummaryV2  [^5] https://github.com/LSSTDESC/OpSimSummaryV2.(yoachim_lsstrubin_sim_2023), that matches simulated observations made with the LSST software OpSim  [^6] https://github.com/lsst/rubin_sim., with galaxy catalogs from Uchuu mocks. In this work we use the output baseline_v3.3_10years of OpSim. We focus on the Wide Fast Deep (WFD) program of the LSST survey that will observe in six filters over a large area of the southern sky. To build our simulation input files, OpSimSummaryV2 represents the LSST WFD survey as a collection of equal-area pixels computed using the HEALPix [^7] http://healpix.sf.net. (gorski_healpix_2005) scheme as implemented in the python library healpy  [^8] https://github.com/healpy/healpy. (zonca_healpy_2019). In this work, we use pixels with an area of $\sim 0.05\text{ deg}^{2}$ and cut pixels with less than 500 observations over the 10 years of the survey. Then, OpSimSummaryV2 associates to each of the pixels the LSST observations and Uchuu UniverseMachine host galaxies that lie within the radius of the Rubin Observatory field-of-view. These pixels are shown in Fig. 2. The holes in the footprint correspond to the LSST Deep Drilling Field program that observes a small area with long exposures at a high cadence. These pixels were removed by imposing the cut that a pixel should be associated with less than 1100 observations. We note that we still have observations at these coordinates in our simulations, since observations are taken within the LSST field radius around each pixel. Finally, 50 000 pixels are randomly drawn and their associated observations and hosts are taken to build the SNANA simulation inputs. Fig. 3 shows the angular distribution of the simulated SNe Ia.
@@
To simulate a realistic number of SNe Ia  we use the SN Ia rate value from frohmaier_volumetric_2019, which is $r_{v,\mathrm{F19}}=2.27\times(1+z)^{1.7}\times 10^{-5}$ SN Ia/Mpc/year and rescale it to our fiducial value of $h$ using
@@
The simulated photometry is obtained by the integration of a spectral template of each SN Ia. In this work, we use the spectral template from the SALT3 model (guy_salt2_2007; kenworthy_salt3_2021) as trained in vincenzi_dark_2024. The SALT3 model gives a spectral template for SNe Ia that depends on three parameters: the magnitude of the SN Ia $m_{b}$, the stretch $x_{1}$, and the color $c$. The absolute magnitude of each SN Ia is defined by following the Tripp equation (tripp_two-parameter_1998) as
@@
where $\mu(\boldsymbol{z_{\mathrm{cos}}})$ is the distance moduli at the redshift of the SN Ia host and the last term is the relativistic beaming from peculiar velocity (davis_effect_2011). In Eqs. 3 and 4 we put in bold the quantities that differ for each SN Ia.
We generate simulations using four different SN Ia intrinsic scatter models. The first is the least realistic and is used for comparison: coherent (achromatic) Gaussian scatter with a standard deviation of $\sigma_{M}=0.12$. We refer to this simulation as $S_{\mathrm{COH}}$. The next two are models that have been used historically in SN Ia analyses: the G10 model (guy_supernova_2010) where 70% of the luminosity scatter is due to achromatic variations and 30% to chromatic ones, and the C11 model (chotard_reddening_2011) where 25% of the scattering is achromatic and 70% is chromatic. These two models directly add a wavelength-dependent scatter to the spectral model provided by SALT3. We refer to these two simulations as $S_{\mathrm{G10}}$ and $S_{\mathrm{C11}}$, respectively. Recent studies, however, have disfavoured these two models (vincenzi_dark_2024). A more realistic model is presented in brout_its_2021 (hereafter referred as BS21), which proposes that the scattering is caused by variation in dust extinction parameters of SN Ia host galaxies. We simulate the brout_its_2021 model using the dust parameters fitted in popovic_pantheon_2023 (hereafter referred as P23). We refer to this simulation as $S_{\mathrm{P23}}$.
@@
with a value for $\gamma$ of 0.05 mag. The distribution of $x_{1}$ and $c$ are also chosen to be identical for these models. We use the parent populations $\mathcal{P}(x_{1}|M_{\mathrm{host}})$ and $\mathcal{P}(c|M_{\mathrm{host}})$ derived in Table 4 and 8 of popovic_improved_2021 (hereafter referred as P21) which aim to model the correlations between SN Ia parameters and host masses.
In the brout_its_2021 model, the Tripp relation of Eq. 3 is slightly modified such as
@@
where $\alpha_{\mathrm{fid}}$ is again fixed at a value of 0.15 and $\mathcal{P}(x_{1}|M_{\mathrm{host}})$ is described by the same parent population as for the $S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$ and $S_{\mathrm{C11}}$ simulations. The brout_its_2021 model assumes that the SN color can be divided into two components. The first corresponds to an intrinsic color variation, $\beta c$, where $\beta$ is a random variable following a Gaussian law $\mathcal{P}(\beta)\sim\mathcal{N}(\mu_{\beta},\sigma_{\beta})$ of parameters
@@
[…]
These values were derived using the method presented in popovic_pantheon_2023 and are presented in Table 3 of vincenzi_dark_2024. We summarize the parameters for our four intrinsic scatter models in Table 2.